# CAD-VAE: Leveraging Correlation-Aware Latents for Comprehensive Fair Disentanglement

### v1

@@ Abstract
While deep generative models have significantly advanced representation learning, they may inherit or amplify biases and fairness issues by encoding sensitive attributes alongside predictive features. Enforcing strict independence in disentanglement is often unrealistic when target and sensitive factors are naturally correlated. To address this challenge, we propose CAD-VAE (Correlation-Aware Disentangled VAE), which introduces a correlated latent code to capture the shared information between target and sensitive attributes. Given this correlated latent, our method effectively separates overlapping factors without extra domain knowledge by directly minimizing the conditional mutual information between target and sensitive codes. A relevance-driven optimization strategy refines the correlated code by efficiently capturing essential correlated features and eliminating redundancy. Extensive experiments on benchmark datasets demonstrate that CAD-VAE produces fairer representations, realistic counterfactuals, and improved fairness-aware image editing. Code
Figure 1: Generated Fair Fine-Grained Image Editing samples from CAD-VAE. The blue-framed section interpolates $z_{Y}$ and $z_{S}$ while fixing $z_{R}$ and $z_{X}$. The red-framed section interpolates $z_{Y}$ and $z_{R}$, replacing $z_{S}$ with the reference while keeping $z_{X}$ constant. $z_{R}$ enables fine control over sensitive relevant features like makeup or mustache. See details in Section 4.3
1 Introduction
@@
Deep generative models have achieved remarkable success in capturing complex data distributions for applications ranging from image synthesis [1] to video generation [34]. In particular, variational autoencoders (VAEs)[13, 21, 31, 33] have provided a principled approach to representation learning, where data are encoded into compact latent variables that effectively capture meaningful factors of variation. However, while these latent representations have enabled impressive performance in numerous tasks, concerns about fairness have emerged, as models can inadvertently learn and amplify biases present in training data[15].
@@
Such fairness issues arise when the target label and sensitive label become entangled due to societal or dataset biases [24, 26]. To address these problems, existing methods commonly fall into two categories. Invariant learning techniques aim to remove sensitive attributes from the learned representation, often via adversarial training or additional regularization [24, 40]. By contrast, disentanglement approaches encourage the model to partition its latent space into separate codes for target and sensitive information, seeking statistical independence among them [6, 28]. Although these solutions have made progress, they typically assume minimal correlation between target and sensitive factors or enforce strict separation via mutual information penalties [3]. However, multiple works [15, 36] have demonstrated that achieving fully fair disentanglement is fundamentally impossible under realistic conditions. First, many datasets contain unwanted correlations between the target label and sensitive attributes due to societal bias, making it infeasible to preserve all predictive cues while completely discarding sensitive information [8]. Second, certain features inherently influence both target and sensitive attributes, so perfectly partitioning features into disjoint latent spaces is unachievable without compromising prediction accuracy [22]. In these circumstances, any attempt at full disentanglement faces an inevitable trade-off between fairness and utility.
@@
A natural way to handle this correlation is to explicitly model how target and sensitive attributes overlap. For instance, some methods rely on causal graphs to separate task-relevant features and capture their relationships with sensitive variables [20, 41, 51, 14]. However, constructing such graphs requires extensive domain knowledge, which is often challenging to acquire in real-world scenarios.
Motivated by the limitations of existing methods, we introduce an additional latent code specifically designed to capture the shared information between target and sensitive attributes. Our approach directly minimizes the conditional mutual information between target and sensitive property predictions, conditioned on this correlated latent code. This design guarantees that the target and sensitive latent codes remain independent once the overlapping factors are absorbed. Moreover, an explicit relevance learning strategy ensures that the correlated latent code captures only the essential shared information without extra domain knowledge. In summary, our contributions are as follows:
@@
2 Related Work
2.1 Fair Disentanglement Learning
Instead of directly learning sensitive-information-free representations [28, 30, 46], Fair Disentanglement Learning focuses on disentangling these representations into two sets of latent variables, each dedicated to capturing target or sensitive information separately [43]. One of the pioneering works, $\beta$-VAE [13], introduced a framework for semantic feature disentanglement by enforcing a stronger prior constraint on the variational distribution. Extending this idea, FactorVAE [19] proposed minimizing the total correlation (TC) among latent variables to achieve a more refined decomposition of latent space. Specifically, FactorVAE introduces a discriminator-based approach to approximate the TC loss as:
$L_{TC}=KL(q(z)||\prod_{j}q(z_{j}))\approx\mathbb{E}{q_{\phi}(z)}\left[\log\frac{D(z)}{1-D(z)}\right],$ (1)
where the discriminator $D$ aims to distinguish whether the latent code $z$ is sampled from the aggregate posterior or from a combination of in-batch permutations of marginal distributions across each latent dimension $j$. Meanwhile, the encoder is trained to fool the discriminator, promoting independence among latent dimensions. Further advancements on FactorVAE, FairFactorVAE [28] incorporate strategies to obfuscate sensitive information within the latent space, enhancing privacy preservation.
Another notable approach, FFVAE [6], extends the use of total correlation [3] specifically to disentangle the sensitive code with adaptability to the dimension of sensitive information. GVAE [7] employs adversarial training to minimize the leakage of unwanted information in each latent code, offering an alternative method to enforce independence between target and sensitive representations. Meanwhile, ODVAE [42] introduces orthogonal priors for target and sensitive codes, ensuring that these codes maintain minimal overlap. FairDisCo [27] approaches the independence challenge through distance covariance minimization, offering a non-adversarial alternative to traditional methods.
2.2 Correlation-Aware Learning
While disentanglement methods advance different information separations, achieving perfectly independent latent codes remains challenging. The need to recover both sensitive and target information often conflicts with maintaining complete independence, as these types of information are frequently correlated in practice [32, 15]. For instance, in facial attribute recognition on the CelebA dataset [29], ’mustache’ (sensitive relevant) correlates with both gender (sensitive) and attractiveness (target), highlighting the inherent dependency between sensitive and target codes.
To address the challenge of correlated information, Correlation-Aware Learning [20, 41, 51, 14] employs a causal graph approach to separate latent variables based on their correlation with sensitive information. However, constructing such graphs requires domain knowledge and may not always be feasible. Additionally, not accounting for causal relationships between variables can hinder the independence of learned latent codes, leading to practical and computational challenges [15].
To address the inherent correlation between target and sensitive information while reducing dependence on domain knowledge, FADES [15] proposes a novel approach that minimizes conditional mutual information by grouping samples based on different attributes to learn shared information in the sensitive relevant code. Although innovative, this indirect method may not effectively eliminate unwanted information leakage and lacks explicit guidance to ensure the sensitive relevant code captures an appropriate and balanced amount of information. These limitations motivate our approach, which directly minimizes conditional mutual information to achieve complete disentanglement and provides explicit constraints to ensure accurate relevance capture.
2.3 Counterfactual Fairness
@@
Counterfactual fairness (CF) ensures fairness by evaluating whether predictions remain consistent when sensitive attributes are altered while keeping other features unchanged, aiming to mitigate bias through comparisons between factual and counterfactual outcomes [23]. The causal inference technique is widely adopted to generate these counterfactual instances [50, 16, 51, 4, 45], enabling a comparative analysis between actual and hypothetical scenarios.
Graph-based models with predefined intervention and individual variables [20, 25] use causal effect analysis to capture feature interactions for realistic counterfactuals. However, their effectiveness relies on precise domain knowledge; without it, generated counterfactuals may become unrealistic or nonsensical, such as depicting a girl with a mustache.
While CF demonstrates strong theoretical foundations, its dependence on precise domain knowledge complicates model design and may limit practical utility. To address this, our approach introduces a correlated latent code with an explicit relevance learning strategy, allowing the model to learn attribute relationships autonomously and enhance counterfactual fairness without external knowledge.
3 Method
3.1 CAD-VAE
@@
Let $\mathcal{D}=\{(x_{i},y_{i},s_{i})\}_{i=1}^{N}$ denote a dataset consisting of triplets, where $x_{i}$ denotes an input sample (e.g., an image), $y_{i}$ is the label of $x_{i}$ corresponding to target property $Y$, and $s_{i}$ is the label corresponding to the sensitive property $S$. The value range of $Y$ and $S$ is $\mathcal{Y}$ and $\mathcal{S}$, respectively, i.e., $y\in\mathcal{Y}$ and $s\in\mathcal{S}$. The goal is to learn a latent representation that factorizes the information pertinent to $Y$, the information pertinent to $S$, the shared information between $Y$ and $S$, and the background or irrelevant factors. We introduce the following latent representations to capture different information:
@@
Hence, for a single observation, the corresponding latent variable set is $z:=(z_{X},z_{Y},z_{S},z_{R})$. We first present the problem definition, model components, and architecture in Section 3.1, which serve as crucial foundations for the subsequent sections. To achieve correlation-aware disentanglement learning, we propose directly minimizing the conditional mutual information between $z_{Y}$ and $z_{S}$ with respect to their corresponding opposite attributes $S$ and $Y$, conditioned on $z_{R}$. This approach is complemented by an explicit relevance learning strategy that constrains $z_{R}$ to effectively capture shared information between $Y$ and $S$ while avoiding redundant information. Detailed explanations of these strategies are provided in Section 3.3 and Section 3.4, respectively.
To learn such latent code, we employ a Variational Autoencoder (VAE) [13] with an encoder $q_{\phi}(z\mid x)$ and a decoder $p_{\theta}(x\mid z)$ as our backbone. The prior distribution over the latent variables $p_{\theta}(z)$ is typically chosen to be factorized as a standard Gaussian: $p_{\theta}(z)=\mathcal{N}(z\mid 0,I)$. To train the VAE, we minimize the negative evidence lower bound (negative ELBO), defined as
$\min_{\theta,\phi}\Bigl{[}\mathcal{L}_{\text{VAE}}(\theta,\phi)\Bigr{]}\;=\;\mathbb{E}_{q_{\phi}(z\mid x)}\bigl{[}-\log p_{\theta}(x\mid z)\bigr{]}$ (2)
$\;+\;\mathrm{KL}\Bigl{(}q_{\phi}(z\mid x)\,\big{\|}\,p_{\theta}(z)\Bigr{)},$
where $\theta$ represents the parameters of the decoder and $\phi$ represents the parameters of the encoder.
@@
  • Eliminating information leakage (see in Section 3.3)
  • Encouraging $z_{R}$ encapsulate only the correlated information $Y\cap S$ (see in Section 3.4)
@@
Here, we first present the training method and loss function of each classifiers.
@@
  • $f_{s}(z_{S},z_{R})$ predicts $\hat{s}$ from $(z_{S},z_{R})$;
@@
  • $f_{s\_op}(z_{Y})$ attempts to predict $\hat{s}$ from $z_{Y}$.
@@
To measure the information leakage, which is illustrated in Section 3.3,we introduce:
@@
3.2 Conditional Independence
As discussed in the Related Works, correlated information between target attribute $Y$ and sensitive attribute $S$ is pervasive in disentanglement learning tasks. To address this challenge, we introduce an additional latent code $z_{R}$ to capture this correlated information, as described in Section 3.1.
The latent code $z_{R}$ specifically directs the overlapping information between $Y$ and $S$, enabling the independent latent codes $z_{Y}$ and $z_{S}$ to remain free of unwanted correlations while preserving the predictiveness of the model [6, 20].
From a causal perspective, if $Y$ and $S$ are conditionally independent given $z_{R}$ (i.e., $Y\perp S\mid z_{R}$), $z_{R}$ effectively serves as the common cause, promoting the independence of $z_{Y}$ and $z_{S}$. Figure 2 illustrates this information flow, demonstrating how $z_{R}$ manages correlated information while maintaining the independence of the other latent codes.
Our fairness objective is to achieve independence between $z_{Y}$ and $z_{S}$ by enforcing conditional independence between $\hat{Y}$ and $\hat{S}$ given $z_{R}$, where: $\hat{Y}\!=\!f_{y}(z_{Y},z_{R})$ and $\hat{S}\!=\!f_{s}(z_{S},z_{R})$. Additionally, we utilize $z_{X}$ as an auxiliary latent code to filter out irrelevant features, allowing for more controlled counterfactual generation.
We first consider enforcing the proposed disentanglement by directly imposing Conditional Independence (CI), namely
$p_{\theta}(\hat{Y},\hat{S}\mid z_{R})=p_{\theta}(\hat{Y}\mid z_{R})\,p_{\theta}(\hat{S}\mid z_{R}),$ (7)
which implies $\hat{Y}\,\perp\,\hat{S}\;\big{|}\;z_{R}$. Ideally, one could measure how close $\hat{Y}$ and $\hat{S}$ are to being conditionally independent by minimizing the divergence $D\!\bigl{(}p_{\theta}(\hat{Y},\hat{S}\mid z_{R})\,\big{\|}\,p_{\theta}(\hat{Y}\mid z_{R})\,p_{\theta}(\hat{S}\mid z_{R})\bigr{)}$, but this is generally intractable to compute in practice. Instead, we follow standard information-theoretic principles and propose to minimize the Conditional Mutual Information (CMI): $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$, as a tractable surrogate objective. By definition of CMI,
@@ 3.2 Conditional Independence
$\int\!D_{\mathrm{KL}}\!\Bigl{(}p_{\theta}(\hat{Y},\hat{S}\!\mid\!z_{R})\,\big{\|}\,p_{\theta}(\hat{Y}\!\mid\!z_{R})\,p_{\theta}(\hat{S}\!\mid\!z_{R})\Bigr{)}\,\!dP_{z_{R}},$
and thus minimizing $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$ is equivalent to driving the above divergence to zero. See Appendix A for a detailed proof of the equivalence between vanishing conditional mutual information and conditional independence.
From information theory, we know $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})=I_{\phi}(\hat{S};\hat{Y}\mid z_{R})$, thus, when $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$ (and equivalently $I_{\phi}(\hat{S};\hat{Y}\mid z_{R})$) is minimized, any undesired dependence between $\hat{Y}$ and $\hat{S}$ is removed.
3.3 Direct Minimization of Conditional Mutual Information
To achieve CI as (7), we propose directly minimizing:
@@ 3.3 Direct Minimization of Conditional Mutual Information
from the same transformation (see detailed derivation in Appendix B):
@@
where $z_{R}$ is determined here, $H_{\phi}(\hat{Y}\mid z_{R})$ is a constant, henceforth we need to minimize $-H_{\phi}(\hat{Y}\mid z_{S},z_{R})$. Empirically, we directly minimize the lower bound of it: $-H_{\phi}(\hat{Y}\mid z_{S})$, since: $-H_{\phi}(\hat{Y}\mid z_{S},z_{R})\geq-H_{\phi}(\hat{Y}\mid z_{S})$. Symmetrically, the minimization of $I_{\phi}(\hat{S};z_{Y}\mid z_{R})$ shown in (3.3) is the same concept, see Appendix B for detailed derivation. In this optimization process, $z_{R}$ is responsible for containing any correlation information between target attribute $Y$ and sensitive attribute $S$.

### v2

@@ Abstract
While deep generative models have significantly advanced representation learning, they may inherit or amplify biases and fairness issues by encoding sensitive attributes alongside predictive features. Enforcing strict independence in disentanglement is often unrealistic when target and sensitive factors are naturally correlated. To address this challenge, we propose CAD-VAE (Correlation-Aware Disentangled VAE), which introduces a correlated latent code to capture the information shared between the target and sensitive attributes. Given this correlated latent, our method effectively separates overlapping factors without extra domain knowledge by directly minimizing the conditional mutual information between target and sensitive codes. A relevance-driven optimization strategy refines the correlated code by efficiently capturing essential correlated features and eliminating redundancy. Extensive experiments on benchmark datasets demonstrate that CAD-VAE produces fairer representations, realistic counterfactuals, and improved fairness-aware image editing. Source code is available : https://github.com/merry7cherry/CAD-VAE
Introduction
@@
Deep generative models have achieved remarkable success in capturing complex data distributions for applications ranging from image synthesis (Bai et al. 2024; Ma et al. 2025c, b) to video generation (Montanaro et al. 2024). In particular, variational autoencoders (VAEs)(Higgins et al. 2017; Kingma and Welling 2022; Mathieu et al. 2019; Ma et al. 2025a) have provided a principled approach to representation learning, where data are encoded into compact latent variables that effectively capture meaningful factors of variation. However, while these latent representations have enabled impressive performance in numerous tasks, concerns about fairness have emerged, as models can inadvertently learn and amplify biases present in training data(Jang and Wang 2024).
@@
Such fairness issues arise when the target label and sensitive label become entangled due to societal or dataset biases (Lahoti et al. 2020; Liu et al. 2021). To address these problems, existing methods commonly fall into two categories. Invariant learning techniques aim to remove sensitive attributes from the learned representation, often via adversarial training or additional regularization (Lahoti et al. 2020; Roy and Boddeti 2019). By contrast, disentanglement approaches encourage the model to partition its latent space into separate codes for target and sensitive information, seeking statistical independence among them (Creager et al. 2019; Liu, Sun, and Zhao 2023). Although these solutions have made progress, they typically assume minimal correlation between target and sensitive factors or enforce strict separation via mutual information penalties (Chen et al. 2018). However, multiple works (Jang and Wang 2024; Park et al. 2020) have demonstrated that achieving fully fair disentanglement is fundamentally impossible under realistic conditions. First, many datasets contain unwanted correlations between the target label and sensitive attributes due to societal bias, making it infeasible to preserve all predictive cues while completely discarding sensitive information (Dressel and Farid 2018). Second, certain features inherently influence both target and sensitive attributes, so perfectly partitioning features into disjoint latent spaces is unachievable without compromising prediction accuracy (Kohavi 1996). In these circumstances, any attempt at full disentanglement faces an inevitable trade-off between fairness and utility.
@@
A natural way to handle this correlation is to explicitly model how target and sensitive attributes overlap. For instance, some methods rely on causal graphs to separate task-relevant features and capture their relationships with sensitive variables (Kim et al. 2021; Sánchez-Martin, Rateike, and Valera 2022; Zhu et al. 2023; Hwa et al. 2024). However, constructing such graphs requires extensive domain knowledge, which is often challenging to acquire in real-world scenarios.
Motivated by the limitations of existing methods, an correlated latent code is introduced to capture the shared information between target and sensitive attributes. Our approach advances existing methods (Jang and Wang 2024; Park et al. 2020) by a directly minimizing conditional mutual information mechanism to achieve disentanglement and an explicit relevance learning strategy to learn the correlated latent code efficiently and properly, as summarized follow:
@@
Related Work
Fair Disentanglement Learning
Fair disentanglement methods aim to separate representations into target-related and sensitive-related latent codes rather than directly removing sensitive information (Liu, Sun, and Zhao 2023; Madras et al. 2018; Xu et al. 2018; Wang et al. 2024). Early works such as $\beta$-VAE (Higgins et al. 2017) and FactorVAE (Kim and Mnih 2018) introduced mechanisms for semantic decomposition of latent factors, with FactorVAE promoting independence across dimensions by reducing total correlation. Building on this foundation, FairFactorVAE (Liu, Sun, and Zhao 2023) further restricts sensitive leakage within the disentanglement process.
Subsequent studies refine these ideas by emphasizing flexible or guided decomposition. FFVAE (Creager et al. 2019) adapts latent structures to better isolate sensitive attributes, while GVAE (Ding et al. 2020) employs adversarial constraints to suppress unwanted information. Other strategies incorporate structural priors such as orthogonality in ODVAE (Sarhan et al. 2020) or distance-covariance minimization in FairDisCo (Liu et al. 2022). These approaches collectively illustrate progress in disentanglement, yet also reveal the challenge of fully separating target and sensitive information when these factors are inherently correlated.
Correlation-Aware Learning
Despite advancements in disentanglement, perfect independence between latent codes is difficult to achieve due to natural correlations between sensitive and target attributes (Mehrabi et al. 2021; Jang and Wang 2024). For example, facial attributes in CelebA (Liu et al. 2015), such as “mustache,” correlate with both gender and attractiveness, complicating clean separation.
Correlation-aware learning frameworks seek to address this by leveraging causal graphs to categorize latent variables according to their relationships with sensitive attributes (Kim et al. 2021; Sánchez-Martin, Rateike, and Valera 2022; Zhu et al. 2023; Hwa et al. 2024). However, causal-graph construction requires strong domain knowledge, and inaccurate assumptions may hinder independence (Jang and Wang 2024). FADES (Jang and Wang 2024) mitigates this dependency by grouping samples across attributes to approximate conditional mutual information and capture shared sensitive-relevant structure. While effective in some contexts, this indirect method may still allow leakage and offers limited control over relevance allocation. These limitations motivate approaches that directly optimize conditional independence with explicit guidance for balancing sensitive-relevant information.
Counterfactual Fairness
@@
Counterfactual fairness (CF) evaluates whether predictions remain stable when sensitive attributes are hypothetically altered (Kusner et al. 2017). Causal inference is widely used to generate counterfactual instances (Zhou et al. 2024; Jung et al. 2025; Zhu et al. 2023; Chiappa 2019; Wu, Zhang, and Wu 2019), enabling comparisons between factual and hypothetical outcomes.
Graph-based CF models (Kim et al. 2021; Li et al. 2025) rely on predefined causal structures to produce realistic counterfactuals, but these structures demand precise domain knowledge; inaccurate models may lead to implausible counterfactuals, such as depicting a female subject with a mustache. Although CF offers strong theoretical grounding, its reliance on domain expertise limits practical deployment. To overcome this challenge, our approach introduces a correlated latent code with an explicit relevance-learning mechanism, allowing the model to autonomously learn attribute relationships and enhance counterfactual fairness without external causal assumptions.
Preliminary
Conditional Independence and Mutual Information
Proposition 1 (Conditional Independence).
Let $A$, $B$, and $C$ be random variables. We say that $A$ is conditionally independent of $B$ given $C$, denoted $A\perp B\mid C$, if and only if their conditional joint probability distribution factorizes as follows:
$p(A,B\mid C)=p(A\mid C)p(B\mid C).$ (1)
Directly measuring the degree of conditional independence by computing the divergence between the two sides of Eq. (1) is often intractable in practice, especially in the context of deep learning models where the underlying distributions are complex and high-dimensional. Instead, a common and more tractable approach is to use an information-theoretic surrogate measure.
Definition 1 (Conditional Mutual Information).
The Conditional Mutual Information (CMI) between two random variables $A$ and $B$ given a third random variable $C$ measures the expected amount of information that $A$ and $B$ share, conditioned on $C$. It is defined as the expected Kullback-Leibler (KL) divergence between the conditional joint distribution and the product of the conditional marginal distributions:
$I(A;B\!\mid\!C)\!=\!$ (2)
$\mathbb{E}_{p(C)}\!\left[D_{\mathrm{KL}}\!\Bigl(p(A,B\!\mid\!C=c)\!\,\big\|\,\!p(A\!\mid\!C=c)p(B\!\mid\!C=c)\Bigr)\right].$
This can be expressed over the entire distribution of $C$ as:
$I(A;B\!\mid\!C)\!=\!\!\int\!D_{\mathrm{KL}}\!\Bigl(p(A,B\!\mid\!C)\!\,\big\|\,\!p(A\!\mid\!C)p(B\!\mid\!C)\!\Bigr)d{p(C)}.$ (3)
CMI provides a principled way to measure conditional dependence due to its fundamental properties.
Lemma 1 (Properties of CMI).
Conditional mutual information is non-negative, i.e., $I(A;B\mid C)\geq 0$. Furthermore, $I(A;B\mid C)=0$ if and only if $A$ and $B$ are conditionally independent given $C$ ($A\perp B\mid C$). See Appendix 1 for a detailed proof.
Lemma 2 (Symmetry of CMI).
Conditional mutual information is symmetric in its primary arguments:
$I(A;B\mid C)=I(B;A\mid C).$ (4)
Variational Autoencoder
A Variational Autoencoder (VAE) (Kingma and Welling 2022) is a generative model that learns a latent representation of data. It uses an encoder network, $q_{\phi}(z\mid x)$, to map an input sample $x$ to a latent distribution, and a decoder network, $p_{\theta}(x\mid z)$, to reconstruct the input from a latent sample $z$.
The model is trained by minimizing the negative Evidence Lower Bound (ELBO):
$\mathcal{L}_{\text{VAE}}(\theta,\phi)\;=\;$ (5)
$\underbrace{\mathbb{E}_{q_{\phi}(z\mid x)}\bigl[-\log p_{\theta}(x\mid z)\bigr]}_{\text{Reconstruction Loss}}\;+\;\underbrace{\mathrm{KL}\Bigl(q_{\phi}(z\mid x)\,\big\|\,p_{\theta}(z)\Bigr)}_{\text{KL Divergence}},$
where the first term measures reconstruction accuracy and the second term is a regularizer that pushes the learned latent distribution $q_{\phi}(z\mid x)$ towards a prior $p_{\theta}(z)$, which is typically a standard Gaussian $\mathcal{N}(0,I)$.
Total Correlation Loss
To enforce statistical independence among latent variables, FactorVAE (Kim and Mnih 2018) introduces a penalty on the Total Correlation (TC). TC is the Kullback-Leibler (KL) divergence between the aggregate posterior, $q(z)$, and the product of its marginals, $\prod_{j}q(z_{j})$:
$L_{TC}=\text{KL}\left(q(z)\middle\|\prod_{j}q(z_{j})\right)$ (6)
As this term is intractable to compute directly, it is approximated using a discriminator, $D$, which is trained to distinguish between samples from $q(z)$ and samples from the product of marginals (approximated by permuting dimensions across a batch). The encoder, in turn, is trained to minimize the following adversarial loss, thereby fooling the discriminator and reducing the TC:
$L_{TC}\approx\mathbb{E}_{q(z)}\left[\log\frac{D(z)}{1-D(z)}\right]$ (7)
Method
We first present the problem definition, model components, and architecture, which serve as crucial foundations for the subsequent sections.
CAD-VAE
@@
Let $\mathcal{D}=\{(x_{i},y_{i},s_{i})\}_{i=1}^{N}$ denote a dataset consisting of triplets, where $x_{i}$ denotes an input sample (e.g., an image), $y_{i}$ is the label of $x_{i}$ corresponding to target property $Y$, and $s_{i}$ is the label corresponding to the sensitive property $S$. The value range of $Y$ and $S$ is $\mathcal{Y}$ and $\mathcal{S}$, respectively, i.e., $y\in\mathcal{Y}$ and $s\in\mathcal{S}$.
As discussed in Introduction, correlated information between the target attribute $Y$ and the sensitive attribute $S$ is pervasive in disentanglement learning. To address this, we introduce an additional latent code $z_{R}$ to explicitly model this correlated information.
The goal is to learn a latent representation that factorizes the information relevant to $Y$, the information relevant to $S$, the shared information between $Y$ and $S$, and the background or irrelevant factors. As defined below:
@@
Hence, for a single observation, the corresponding latent variable set is $z:=(z_{X},z_{Y},z_{S},z_{R})$.
The latent code $z_{R}$ isolates the overlapping information between $Y$ and $S$, which allows the primary latent codes $z_{Y}$ and $z_{S}$ to remain free of unwanted correlations while preserving the model’s predictive power (Creager et al. 2019; Kim et al. 2021; Jang and Wang 2024). From a causal perspective as illustrated in Figure 1, if $Y$ and $S$ are conditionally independent given $z_{R}$, then $z_{R}$ acts as their common cause, thereby promoting the independence of $z_{Y}$ and $z_{S}$.
To learn such latent code, we employ the Variational Autoencoder (VAE) framework as our backbone. The model is trained to minimize the negative ELBO, $\mathcal{L}_{\text{VAE}}$, as defined in Eq. (LABEL:vae_loss).
@@
  • Eliminating information leakage (see in subsequent section)
  • Encouraging $z_{R}$ encapsulate only the correlated information $Y\cap S$ (see in subsequent section)
Here, we first present the training method and loss function of each classifier.
@@
  • $f_{s}(z_{S},z_{R})$ is a classifier that predicts $\hat{s}$ from $(z_{S},z_{R})$;
@@
  • $f_{s\_op}(z_{Y})$ is an opponent classifier that attempts to predict $\hat{s}$ from $z_{Y}$.
@@
To measure the information leakage, we introduce:
@@
To achieve correlation-aware disentanglement learning, we propose directly minimizing the conditional mutual information between $z_{Y}$ and $z_{S}$ with respect to their corresponding opposite attributes $S$ and $Y$, conditioned on $z_{R}$. This approach is complemented by an explicit relevance learning strategy that constrains $z_{R}$ to effectively capture shared information between $Y$ and $S$ while avoiding redundant information. Detailed explanations of these strategies are provided in following section.
Conditional Independence for Disentanglement
Our fairness objective is to achieve independence between $z_{Y}$ and $z_{S}$ by enforcing conditional independence between their respective predictions, $\hat{Y}$ and $\hat{S}$, given the shared latent code $z_{R}$. The predictions are generated by classifiers: $\hat{Y}=f_{y}(z_{Y},z_{R})$ and $\hat{S}=f_{s}(z_{S},z_{R})$. This objective is formally expressed as $\hat{Y}\perp\hat{S}\mid z_{R}$. Following Proposition 1, this conditional independence is equivalent to the factorization of the conditional joint distribution:
$p_{\theta}(\hat{Y},\hat{S}\mid z_{R})=p_{\theta}(\hat{Y}\mid z_{R})p_{\theta}(\hat{S}\mid z_{R}).$ (12)
As noted in the Preliminary section, directly minimizing the divergence between the distributions in Eq. (12) is generally intractable. We therefore adopt an information-theoretic approach and minimize the Conditional Mutual Information (CMI), $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$, as a tractable surrogate objective.
From Definition 1, the CMI is the expected KL divergence over $z_{R}$:
@@
$\!\int\!\!D_{\mathrm{KL}}\!\Bigl(p_{\theta}(\hat{Y},\hat{S}\!\mid\!z_{R})\!\,\big\|\,\!p_{\theta}(\hat{Y}\!\mid\!z_{R})\!\,p_{\theta}(\hat{S}\!\mid\!z_{R})\!\Bigr)dP_{z_{R}}.$
According to Lemma 1, minimizing $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$ to zero is equivalent to enforcing the conditional independence defined in Eq. (12). Furthermore, leveraging the symmetry property from Lemma 2, we note that $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})=I_{\phi}(\hat{S};\hat{Y}\mid z_{R})$. Thus, minimizing this CMI ensures that any undesired dependence between $\hat{Y}$ and $\hat{S}$ not explained by $z_{R}$ is removed.
Direct Minimization of Conditional Mutual Information
While FADES (Jang and Wang 2024) minimizing CMI (13) through approximation $I_{\phi}(\hat{Y};S\mid z_{R})$:
$\min_{\phi}\Bigl[I_{\phi}(\hat{Y};S\mid z_{R})\Bigr]=\min_{\phi}\Bigl[H_{\phi}(\hat{Y}\mid z_{R})-H_{\phi}(\hat{Y}\mid S,z_{R})\Bigr]$
by reducing CMI through ground truth-based sample grouping, its reliance on batch-level sampling introduces instability.
In contrast, our method directly minimizes CMI (13) via a principled information-theoretic approach, providing a more robust and stable disentanglement process by avoiding sampling variance and reducing dependency on batch-specific dynamics. To achieve CI as (12), we propose directly minimizing:
@@
from the same transformation (see detailed derivation in Appendix 2):
@@
where $z_{R}$ is determined here, $H_{\phi}(\hat{Y}\mid z_{R})$ is a constant, henceforth we need to minimize $H_{\phi}(\hat{Y}\mid z_{S},z_{R})$. Empirically, we directly minimize the lower bound of it: $H_{\phi}(\hat{Y}\mid z_{S})$, since: $H_{\phi}(\hat{Y}\mid z_{S},z_{R})\geq-H_{\phi}(\hat{Y}\mid z_{S})$. Symmetrically, the minimization of $I_{\phi}(\hat{S};z_{Y}\mid z_{R})$ shown in (Direct Minimization of Conditional Mutual Information) is the same concept, see Appendix 2 for detailed derivation. In this optimization process, $z_{R}$ is responsible for containing any correlation information between target attribute $Y$ and sensitive attribute $S$.